OpenStudy (anonymous):

Use basic identities and algebra to simplify the expression. Assume all denominators are nonzero. (sin t - cost t)^2

myininaya (myininaya):

(sint-cost)(sint-cost)=(sint)^2-2sintcost+(cost)^2

myininaya (myininaya):

but (sint)^2+(cost)^2=1

myininaya (myininaya):

1-2sintcost

myininaya (myininaya):

=1-sin(2x)

OpenStudy (anonymous):

how did you get the last step?